OpenStudy (anonymous):

25^(1/2)=5 convert to logarithmic equation log(25)___=___

OpenStudy (anonymous):

log(25)=2 x log(5)

OpenStudy (anonymous):

when you log something it brings down the exponent such that 1/2 x log (25) =log (5)

OpenStudy (anonymous):

\[b^x=y\iff \log_b(y)=x\]

OpenStudy (anonymous):

multiply both sides by 2 and you get log (25)=2 log(5)

OpenStudy (anonymous):

i think you are supposed to fill in the blanks as follows \[\log_{25}(5)=\frac{1}{2}\]

OpenStudy (anonymous):

that's the format I was looking for. thanks you both:)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

btw if you do these often you should get used to going from \[b^x=y\] to \[\log_b(y)=x\] quickly and with ease

OpenStudy (anonymous):

great! that's useful for sure

OpenStudy (anonymous):

I'll need to practice that..

OpenStudy (anonymous):

\[log_b(a) = k \iff b^k = a\] The \(\iff\) symbol means you can go back and forth from the left side to the right. Each side means the same thing as the other.

OpenStudy (anonymous):

basically the same as what an equals sign says, no?

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

You cannot say that \[log_b(a) = k = b^k = a\] Because \(log_b(a) \) doesn't equal \(a\)

OpenStudy (anonymous):

okay, I'm beginning to see this...

OpenStudy (anonymous):

\[2^4=16\iff \log_2(16)=4\] \[\log_{10}(1000000)=6\iff 10^x=1000000\] \[\log(.01)=-2\iff 10^{-2}=.01\] \[8^{\frac{2}{3}}=4\iff \log_8(4)=\frac{2}{3}\]